OpenStudy (anonymous):

If limx> a[f(x)+g(x)] = 2 and lim x>a [f(x)-g(x)] = 1. Find lim x>a [f(x)g(x)].

OpenStudy (calculusfunctions):

Do you know the properties of limits?

OpenStudy (anonymous):

Yes. But I'm confused what to do with this question.

OpenStudy (calculusfunctions):

\[\lim_{x \rightarrow a}[f(x)+g(x)]=\lim_{x \rightarrow a}f(x)+\lim_{x \rightarrow a}g(x)\] \[\lim_{x \rightarrow a}[f(x)-g(x)]=\lim_{x \rightarrow a}f(x)-\lim_{x \rightarrow a}g(x)\] \[\lim_{x \rightarrow a}[f(x)g(x)]=\lim_{x \rightarrow a}f(x)∙\lim_{x \rightarrow a}g(x)\] There are of course more properties but these are the three you need to solve this problem.

OpenStudy (calculusfunctions):

Do you want a hint?

OpenStudy (anonymous):

Yes please!

OpenStudy (calculusfunctions):

Sorry about that. I was typing and the stupid server keeps resetting. Give me a second to type it again.

OpenStudy (calculusfunctions):

Hint: Let\[\lim_{x \rightarrow a}f(x)=m\]and\[\lim_{x \rightarrow a}g(x)=n\]Do you understand what to do next?

OpenStudy (anonymous):

I think so.... i will retry the question. Thank you very much!

OpenStudy (calculusfunctions):

I'll wait. Show me your work.

OpenStudy (calculusfunctions):

Can you show me the next step using this hint and the properties of limits?

OpenStudy (anonymous):

\[\lim_{x \rightarrow a} f(x) = \lim_{x \rightarrow a} (\frac{ 1 }{ 2 } [f(x)+g(x)] - [\frac{ 1 }{ 2 } f(x)-g(x)]\]

OpenStudy (anonymous):

sorry the minus between the two functions is suppose to be a plus*

OpenStudy (calculusfunctions):

No, I meant that\[\lim_{x \rightarrow a}f(x)+\lim_{x \rightarrow a}g(x)=2\]m + n = 2 (equation 1)\[\lim_{x \rightarrow a}f(x)-\lim_{x \rightarrow a}g(x)=1\]m - n = 1 (equation 2) Now solve for m and n first. What do you get?

OpenStudy (anonymous):

n=1/2 & m=5/2*

OpenStudy (calculusfunctions):

Yes m = 1/2 but n is incorrect. Please try again.

OpenStudy (calculusfunctions):

Sorry for a delay in responding but some idiot keeps harassing by sending me improper messages.

OpenStudy (anonymous):

It's okay. Is there no option to block them?

OpenStudy (calculusfunctions):

I reported him so let's see if they do anything about it.

OpenStudy (anonymous):

is m = 3/2?

OpenStudy (calculusfunctions):

Yes sorry I meant m = 3/2 and n = 1/2. Sorry I'm trying to focus but this person just doesn't stop.

OpenStudy (anonymous):

Beset thing to do is not reply to him cause you're just feeding him to go on and on.

OpenStudy (calculusfunctions):

You're absolutely right.

OpenStudy (calculusfunctions):

So do you know the next step?

OpenStudy (anonymous):

would you use the third limit property (from above) and multiply the two?

OpenStudy (calculusfunctions):

Actually the final answer now that you have solved the individual limits.

OpenStudy (calculusfunctions):

Yes! Go ahead.

OpenStudy (calculusfunctions):

Therefore the product of these limits is?

OpenStudy (anonymous):

3/4

OpenStudy (calculusfunctions):

Beautiful!! Excellent work!! So do you now understand the concept of solving these types of problems?

OpenStudy (anonymous):

Yes, I do. thank you very much!! My prof explained it very complicated... so I didn't really understand it. Really appreciate the help!

OpenStudy (calculusfunctions):

Absolutely any time! I have to log out now. That's what I said after helping the last person as well LOL but I really do have to go. I'll be back in a few hours maybe, so if you need me thensimply message me and I will definitely help you as soon I get back. Take care!

OpenStudy (anonymous):

LOL don't worry that happens to everyone. Thanks once again!!! (:

OpenStudy (calculusfunctions):

Welcome! Bye for now!